OpenStudy (anonymous):

Can someone plzz help im stumped 4. In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x)

OpenStudy (anonymous):

Complete the table of values. x 100 + 23x 90(1.2x) 0 1 2 3 4 5 Use the table to determine at approximately which point the number of cells will be the same for each organism. Graph the system of equations and show the point of intersection. Explain what the points graphed for each line represent. Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c. Find the point(s) of intersection, and explain what the intersection represents in the context of the problem.

OpenStudy (anonymous):

you basically have to substitute for x values given into the equation, for example for 100+23x , when x=0, you write 0 in place of the x, so it would be 100+23*0 = 100 , you have to do the same for the rest of the x values for both functions

OpenStudy (anonymous):

After that, you will notice that for a particular value of x, you get the same end result for each function, this will be your solution to the equation 100+23x = 90(1.2x) , and can be seen as a point of intersection for both lines representing their particular functions